# Whirling injections, surjections, and other functions between finite sets

### v3

@@ 2.5. Consequences of the homomesy
$\mathcal{P}_{r}(n)=\left\{(v_{1},v_{2},\ldots,v_{n})\in[0,1]^{n}\;|\;r\leq v_{1}+v_{2}+\cdots+v_{n}\leq n-r\right\}.$
@@
Let $n=7,r=2$, and consider the vector $(0.9,0.6,0.5,0,0,1,0.4)\in\mathcal{P}_{2}(7)$. We will apply the toggles $T=t_{7}\circ\cdots t_{2}\circ t_{1}$ in a left-to-right order. To be in $\mathcal{P}_{2}(7)$, the sum of the entries must be between 2 and 5. We first toggle $t_{1}$. The first entry is $0.9$, and this entry can be anything from 0 to 1 and still yield of vector in $\mathcal{P}_{2}(7)$. So $t_{1}$ changes the first entry to $0+1-0.9=0.1$. Now we have $(0.1,0.6,0.5,0,0,1,0.4)$. To do the toggle $t_{2}$, the second entry can be anything from 0 to 1 and still be in $\mathcal{P}_{2}(7)$, so we change it from $0.6$ to $0+1-0.6=0.4$, and we now have $(0.1,0.4,0.5,0,0,1,0.4)$. We now do the toggle $t_{3}$. The third entry is $0.5$, but the sum of the other entries is $1.9$, so the third entry must lie in the interval $[0.1,1]$. Thus $t_{3}$ changes the third entry to $0.1+1-0.5=0.6$. Applying $t_{4},t_{5},t_{6},t_{7}$, we get that
@@
We have tested several random orbits and we conjecture that the homomesy in Corollary 2.25 extends to $\mathcal{P}_{r}(n)$. There is one caveat though. A vector with irrational entries probably does not always yield a finite orbit. Thus, we refer to the notion of asympotic homomesy, discussed in [Vor18, Jos19, DL24], and which is clearly equivalent to the previous definition of homomesy over a finite orbit.
@@
Suppose we have a set $\mathcal{S}$, a map $w:\mathcal{S}\rightarrow\mathcal{S}$, and a function (“statistic”) $f:\mathcal{S}\rightarrow\mathbb{R}$. If there exists $c\in\mathbb{R}$ such that for every $x\in\mathcal{S}$,
@@
3. Whirling restricted growth words
A partition of the set $[n]$ is an unordered collection of disjoint subsets of $[n]$, called blocks, whose union is $[n]$. For example $\{\{1,2,5\},\{4,7\},\{3,6\}\}$ is a partition of $[7]$. We often write a partition by writing each block without set braces or commas, and using a bar to separate blocks. For example, $\{\{1,2,5\},\{4,7\},\{3,6\}\}$ would be written $125|47|36$. Notice that $125|47|36=125|36|47=63|152|47$ because the order of the blocks is unimportant as well as the order of the numbers within each block.
@@ 3. Whirling restricted growth words
Let $\operatorname{RG}(n,k)$ denote the set of RG-words corresponding to partitions of $[n]$ with exactly $k$ blocks and $\operatorname{RG}(n)$ the set of RG-words corresponding to all partitions of $[n]$ (without specifying the number of blocks). That is
$\operatorname{RG}(n)=\bigcup\limits_{k=1}^{n}\operatorname{RG}(n,k).$
It is clear that any RG-word corresponds to a unique set partition, and that an RG-word of length $n$ is in $\operatorname{RG}(n,k)$ where $k$ is the maximum entry.
Let $\left\{n\atop k\right\}$ denote the number of partitions of $[n]$ with exactly $k$ blocks and $B_{n}$ denote the total number of partitions of $[n]$. These are called the Stirling numbers of the second kind and Bell numbers respectively; e.g., see [Sta11]. RG-words were first introduced by Hutchinson [Hut63] and have been studied recently by Cai and Readdy [CR17] for their connections to the $q$ and “$q$-$(1+q)$” analogues for Stirling numbers of the second kind.
@@
  • A function $f:[n]\rightarrow[k]$ is in $\operatorname{RG}(n,k)$ if and only if it is surjective and for all $11\leq j\leq k-$ 1, $\min\{i\;|\;f(i)=j\}\leq\min\{i\;|\;f(i)=j+1\}$.
  • A function $f:[n]\rightarrow[n]$ is in $\operatorname{RG}(n)$ if and only if for any $11\leq j\leq n-$ 1,
@@
      $\min\{i\;|\;f(i)=j\}\leq\min\{i\;|\;f(i)=j+1\}$.
The condition $\min\{i\;|\;f(i)=j\}\leq\min\{i\;|\;f(i)=j+1\}$ means that in the one-line notation of $f$, the first occurrence of $j$ occurs before the first $j+1$. This is where the term “restricted growth” comes from.
@@
    $(w_{i}(f))(i)=\left\{\begin{array}[]{ll}j+1&\text{if there exists $i^{\prime}<i$ such that $f(i^{\prime})=j$,}\\ j&\text{if the only $i^{\prime}<\min\{h\;|\;f(h)=j+1\}$ s.t. $f(i^{\prime})=j$is $i^{\prime}=i$,}\\ 1&\text{otherwise.}\end{array}\right.$
@@
    $(w_{i}(f))(i)=\left\{\begin{array}[]{ll}k&\text{if the only value $i^{\prime}$for which $f(i^{\prime})=k$ is $i^{\prime}=i$,}\\ 1&\text{otherwise.}\end{array}\right.$
Refer to Example 3.6 for an example of Proposition 3.4. This says that in the one-line notation of $f$, if the value in position $i$ is $j\not=k$, then $w_{i}$ adds $1$ to it if it is not the first occurrence of $j$. If it is the first occurrence of $j$, then $w_{i}$ leaves it alone if there is not another $j$ to the left of the first $j+1$, and otherwise changes the value to $1$. In the case where the value in position $i$ is $k$, $w_{i}$ leaves it alone if it is the only $k$, and otherwise changes it to $1$.
@@
Case 1a: There exists $i^{\prime}<i$ such that $f(i^{\prime})=j$. Then changing the value of $f(i)$ will not change the RG-word criterion $\min\{i\;|\;f(i)=j\}\leq\min\{i\;|\;f(i)=j+1\}$ from Proposition 3.3. So $(w_{i}(f))(i)=j+1$.
Case 1b: There does not exist $i^{\prime}<i$ such that $f(i^{\prime})=j$. Then we cannot change the value of $f(i)$ to anything larger than $j$ and still have an RG-word. If the only $i^{\prime}<\min\{h\;|\;f(h)=j+1\}$ such that $f(i^{\prime})=j$ is $i^{\prime}=i$, then changing the value of $f(i)$ to something other than $j$ will violate $\min\{i\;|\;f(i)=j\}\leq\min\{i\;|\;f(i)=j+1\}$ and not be an RG-word. Otherwise, changing the value of $f(i)$ to 1 results in an RG-word.
@@
$(w_{i}(f))(i)=\left\{\begin{array}[]{ll}j+1&\text{if there exists $i^{\prime}<i$ such that $f(i^{\prime})=j$,}\\ j&\text{if the only $i^{\prime}<\min\{h\;|\;f(h)=j+1\}$ s.t. $f(i^{\prime})=j$is $i^{\prime}=i$,}\\ 1&\text{otherwise.}\end{array}\right.$
If there is no $h$ satisfying $f(h)=j+1$, then we are not in the second case.
@@
$12213421213341\stackrel{{\scriptstyle w_{1}}}{{\mapsto}}1213341\stackrel{{\scriptstyle w_{2}}}{{\mapsto}}1213341\stackrel{{\scriptstyle w_{3}}}{{\mapsto}}1223341\stackrel{{\scriptstyle w_{4}}}{{\mapsto}}1221341\stackrel{{\scriptstyle w_{5}}}{{\mapsto}}1221341\stackrel{{\scriptstyle w_{6}}}{{\mapsto}}1221341\stackrel{{\scriptstyle w_{7}}}{{\mapsto}}$ 1221342
@@
$12213121213341\stackrel{{\scriptstyle w_{1}}}{{\mapsto}}1213341\stackrel{{\scriptstyle w_{2}}}{{\mapsto}}1213341\stackrel{{\scriptstyle w_{3}}}{{\mapsto}}1223341\stackrel{{\scriptstyle w_{4}}}{{\mapsto}}1221341\stackrel{{\scriptstyle w_{5}}}{{\mapsto}}1221341\stackrel{{\scriptstyle w_{6}}}{{\mapsto}}1221311\stackrel{{\scriptstyle w_{7}}}{{\mapsto}}$ 1221312
@@
Propositions 3.4 and 3.5 show that for $\mathcal{F}=\operatorname{RG}(n,k)$ or $\mathcal{F}=\operatorname{RG}(n)$, whirling at index $i$ either adds 1 to the value $f(i)$ or leaves $f(i)$ alone or changes the value $f(i)$ to 1.
@@
See Figure 6 for an illustration of Theorem 3.8 for the orbits under the action of $\mathbf{w}$ (i.e., $\pi$ is the identity) over $\mathcal{F}=\operatorname{RG}(5,3)$. This theorem is another instance where there is homomesy under an action that produces unpredictable orbit sizes, and for which the order of the map is unknown in general.
@@
Figure 6. The two $\mathbf{w}$-orbits for $\mathcal{F}=\operatorname{RG}(5,3)$. The left orbit has length 16 and the right orbit has length 9. Notice that in the left orbit, there are the same numbers (four each) of functions $h$ satisfying each of $h(2)=1$, $h(3)=1$, $h(4)=1$, and $h(5)=1$. In the right orbit, there are also the same numbers (two each) of functions $h$ satisfying each of $h(2)=1$, $h(3)=1$, $h(4)=1$, and $h(5)=1$. This illustrates Theorem 3.8 for the case $n=5$, $k=3$, and $\pi$ is the identity.
@@
Let $\mathcal{F}$ be either $\operatorname{RG}(n)$ or $\operatorname{RG}(n,k)$ for some $1\leq k\leq n$. Let $\pi,\sigma$ be permutations on $\{2,3,\dots,n\}$ where $\sigma(i)=\pi(i+1)$ for $12\leq i\leq n-$ 1 and $\sigma(n)=\pi(2)$. Then for any $\mathbf{w}_{\pi}$-orbit $\mathcal{O}=(f_{1},f_{2},\dots,f_{\ell})$, there is a $\mathbf{w}_{\sigma}$-orbit $\mathcal{O}^{\prime}=(f_{1}^{\prime},f_{2}^{\prime},\dots,f_{\ell}^{\prime})$ for which $f_{j}^{\prime}=w_{\pi(2)}(f_{j})$ for all $j\in[\ell]$.
@@
For $h\not=\pi(2)$, $f_{i}(h)=f_{i}^{\prime}(h)$. Also, $f_{i}^{\prime}(\pi(2))=f_{i+1}(2)$. Thus for any $h\in[n]$, the multiset of values of $f(h)$ is the same for $f$ ranging over $\mathcal{O}$ is the same as that for $\mathcal{O}^{\prime}$. ∎
@@
For example, on $\mathcal{F}=\operatorname{RG}(6)$, $w_{3}\circ w_{5}\circ w_{4}\circ w_{6}\circ w_{2}$ and $w_{4}\circ w_{6}\circ w_{2}\circ w_{3}\circ w_{5}$ satisfy the above orbit correspondence. However in general two orders for applying the whirling maps do not yield the same orbit structure. For example, $w_{3}\circ w_{5}\circ w_{4}\circ w_{6}\circ w_{2}$ does not have the same orbit structure as $w_{5}\circ w_{6}\circ w_{3}\circ w_{2}\circ w_{4}$ because these are not cyclic rotations of each other.
@@
Let $\mathcal{F}$ be either $\operatorname{RG}(n)$ or $\operatorname{RG}(n,k)$ for some $1\leq k\leq n$ and $\pi$ be a permutation of $\{2,3,\dots,n\}$. Let $i=\pi(2)$ and suppose $12\leq i\leq n-$ 1. Then for $f\in\mathcal{F}$, if $f(i)=1$, then $\left(\mathbf{w}_{\pi}^{-1}(f)\right)(i)\geq\left(\mathbf{w}_{\pi}^{-1}(f)\right)(i+1)$.
@@
Let $f\in\mathcal{F}$ such that $f(i)=1$. Let $g=\mathbf{w}_{\pi}^{-1}(f)$. Since $w_{i}$ is the first map we apply when we apply $\mathbf{w}_{\pi}$ to $g$, $w_{i}$ changes the value of $g(i)$ to 1. By Propositions 3.4 and 3.5, either $i$ is the least $i^{\prime}$ satisfying $g(i^{\prime})=g(i)$, or $g(i)=k$ (the latter option only being possible or making sense if $\mathcal{F}=\operatorname{RG}(n,k)$).
Assume by way of contradiction that $g(i)<g(i+1)$. This could only be possible in the case where $i$ is the least $i^{\prime}$ satisfying $g(i^{\prime})=g(i)$. From the restricted growth condition, $g(i+1)=g(i)+1$. Then Propositions 3.4 and 3.5 say that $w_{i}$ would leave the value $g(i)$ as is, not change it to 1, which is a contradiction. Thus, $g(i)\geq g(i+1)$. ∎
@@
The condition $f(i)=1$ means $1$ and $i$ are in the same block of the partition of $[n]$ corresponding to $f$. Such set partitions can be formed by first choosing a partition of $\{2,3,\dots,n-1\}$ and placing 1 in the same block as $i$. ∎
@@
Let $\mathcal{O}$ be a $\mathbf{w}_{\pi}$-orbit. The goal is to prove that in $\mathcal{O}$, there are the same number of functions $f\in\mathcal{O}$ satisfying $f(i)=1$ as there are satisfying $f(j)=1$, when $2\leq i,j\leq n$. We will approach this by showing that $\mathcal{O}$ has the same number of functions $f$ satisfying $f(i)=1$ as $f(i+1)=1$, for $12\leq i\leq n-$ 1.
Let $i\in\{2,3,\dots,n-1\}$. Without loss of generality, assume $i=\pi(2)$. This is because by Remark 3.10, we could cyclically change $\pi$ to get this condition, without changing the property we wish to prove for orbits of $\mathbf{w}_{\pi}$.
@@
By the way whirling at an index works, either $(\mathbf{w}_{\pi}(g))(i)=g(i)+1$ or $(\mathbf{w}_{\pi}(g))(i+1)=1$.
@@
By Lemma 3.11, $\left(\mathbf{w}_{\pi}^{m}(f)\right)(i)\not=1$ for all $1\leq m\leq r$. Therefore, given any function $f\in\mathcal{O}$ satisfying $f(i)=1$ and $f(i+1)\not=1$, as we repeatedly apply $\mathbf{w}_{\pi}$ to $f$, we obtain a function that sends $i+1$ to 1 before we get another function that sends $i$ to 1. This implies #$\#\{f\in\mathcal{O}\;|\;f(i)=1\}\leq\#\{f\in\mathcal{O}\;|\;f(i+1)=1\}$.
By Lemma 3.12, #$\#\{f\in\mathcal{F}\;|\;f(i)=1\}=\#\{f\in\mathcal{F}\;|\;f(i+1)=1\}$. Therefore, since #$\#\{f\in\mathcal{O}\;|\;f(i)=1\}\leq\#\{f\in\mathcal{O}\;|\;f(i+1)=1\}$ for every $\mathbf{w}_{\pi}$-orbit $\mathcal{O}$ over $\mathcal{F}$, we must have #$\#\{f\in\mathcal{O}\;|\;f(i)=1\}=\#\{f\in\mathcal{O}\;|\;f(i+1)=1\}$ for every orbit $\mathcal{O}$. ∎
Theorem 3.8 implies another homomesic statistic only for the action of $\mathbf{w}_{\pi}$ only on $\operatorname{RG}(n,k)$.
@@
Let $n\geq 2$ and $\pi$ a permutation of $\{2,3,\dots,n\}$. Under the action of $\mathbf{w}_{\pi}$ on $\operatorname{RG}(n,k)$, the statistic
@@
is homomesic with average $k(k-2)$.
@@
Suppose instead that there are $c$ functions $f\in\mathcal{O}$ that satisfy $f(2)=1$. Then this decreases the total value of (${k\choose 2}f(2)$ across $\mathcal{O}$ by (${k\choose 2}c$.
@@
Hence this statistic has average $k(k-2)$ on $\mathcal{O}$. ∎
We finish the paper with some conjectural homomesies for whirling on the set of RG-words corresponding to noncrossing partitions.

### v4

@@ 2.7. Lifting whirling to the piecewise-linear realm
$\mathcal{P}_{r}(n):=\left\{(v_{1},v_{2},\ldots,v_{n})\in[0,1]^{n}:r\leq v_{1}+v_{2}+\cdots+v_{n}\leq n-r\right\}.$
@@
Let $n=7,r=2$, and consider the vector $(0.9,0.6,0.5,0,0,1,0.4)\in\mathcal{P}_{2}(7)$. We will apply the toggles $T=t_{7}\circ\cdots t_{2}\circ t_{1}$ in a left-to-right order. To be in $\mathcal{P}_{2}(7)$, the sum of the entries must be between 2 and 5. We first toggle $t_{1}$. The first entry is $0.9$, and this entry can be anything from 0 to 1 and still yield of vector in $\mathcal{P}_{2}(7)$. So $t_{1}$ changes the first entry to 0+10.9=0.10+1-0.9=0.1. Now we have $(0.1,0.6,0.5,0,0,1,0.4)$. To do the toggle $t_{2}$, the second entry can be anything from 0 to 1 and still be in $\mathcal{P}_{2}(7)$, so we change it from $0.6$ to 0+10.6=0.40+1-0.6=0.4, and we now have $(0.1,0.4,0.5,0,0,1,0.4)$. We now do the toggle $t_{3}$. The third entry is $0.5$, but the sum of the other entries is $1.9$, so the third entry must lie in the interval $[0.1,1]$. Thus $t_{3}$ changes the third entry to 0.1+10.5=0.60.1+1-0.5=0.6. Applying $t_{4},t_{5},t_{6},t_{7}$, we get that
@@
We have tested several random orbits and we conjecture that the homomesy in Corollary 2.23 extends to $\mathcal{P}_{r}(n)$. There is one caveat though. A vector with irrational entries probably does not always yield a finite orbit. Thus, we refer to the notion of asympotic homomesy, discussed in [Vor18, Jos19, DL24], and which is clearly equivalent to the previous definition of homomesy when all of the orbits are finite.
@@
Suppose we have a set $\mathcal{S}$, a map $w:\mathcal{S}\rightarrow\mathcal{S}$, and a function (statistic) $f:\mathcal{S}\rightarrow\mathbb{R}$. If there exists $c\in\mathbb{R}$ such that for every $x\in\mathcal{S}$,
@@
3. Whirling on restricted growth words
3.1. Basic definitions, properties, and main result
A partition of the set $[n]$ is an unordered collection of disjoint subsets of $[n]$, called blocks, whose union is $[n]$. For example, $\{\{1,2,5\},\{4,7\},\{3,6\}\}$ is a partition of $[7]$. We often write a partition by writing each block without set braces or commas, and using a bar to separate blocks. For example, $\{\{1,2,5\},\{4,7\},\{3,6\}\}$ would be written $125|47|36$. Notice that $125|47|36=125|36|47=63|152|47$ because the order of the blocks is unimportant as well as the order of the numbers within each block.
@@ 3.1. Basic definitions, properties, and main result
Let $\operatorname{RG}(n,k)$ denote the set of RG-words corresponding to partitions of $[n]$ with exactly $k$ blocks and $\operatorname{RG}(n)$ the set of RG-words corresponding to all partitions of $[n]$ (without specifying the number of blocks).
Let $\left\{n\atop k\right\}$ denote the number of partitions of $[n]$ with exactly $k$ blocks and $B_{n}$ denote the total number of partitions of $[n]$. These are called the Stirling numbers of the second kind and Bell numbers respectively; e.g., see [Sta11, §1.9]. RG-words were first introduced by Hutchinson [Hut63] and have been studied more recently by Cai and Readdy [CR17] for their connections to the $q$ and “qq-(1+q)(1+q)” analogues for Stirling numbers of the second kind.
@@
  • A function $f:[n]\rightarrow[k]$ is in $\operatorname{RG}(n,k)$ if and only if it is surjective and for all $1\leq j\leq k-1$, $\min f^{-1}(j)<\min f^{-1}(j+1)$.
  • A function $f:[n]\rightarrow[n]$ is in $\operatorname{RG}(n)$ if and only if for any $1\leq j\leq n-1$,
@@
      $\min f^{-1}(j)<\min f^{-1}(j+1)$.
@@
[…]
@@
[…]
Refer to Example 3.6 for an example of Proposition 3.4.
@@
Case 1a: There exists $i^{\prime}<i$ such that $f(i^{\prime})=j$. Then changing the value of $f(i)$ will not change the RG-word criterion $\min f^{-1}(j)<\min f^{-1}(j+1)$ from Proposition 3.3. So $(w_{i}(f))(i)=j+1$.
Case 1b: There does not exist $i^{\prime}<i$ such that $f(i^{\prime})=j$. Then we cannot change the value of $f(i)$ to anything larger than $j$ and still have an RG-word. If the only $i^{\prime}<\min f^{-1}(j+1)$ such that $f(i^{\prime})=j$ is $i^{\prime}=i$, then changing the value of $f(i)$ to something other than $j$ will violate $\min f^{-1}(j)<\min f^{-1}(j+1)$ and not be an RG-word. Otherwise, changing the value of $f(i)$ to 1 results in an RG-word.
@@
[…]
@@
$1213341\stackrel{{\scriptstyle w_{1}}}{{\mapsto}}1213341\stackrel{{\scriptstyle w_{2}}}{{\mapsto}}1213341\stackrel{{\scriptstyle w_{3}}}{{\mapsto}}1223341\stackrel{{\scriptstyle w_{4}}}{{\mapsto}}1221341\stackrel{{\scriptstyle w_{5}}}{{\mapsto}}1221341\stackrel{{\scriptstyle w_{6}}}{{\mapsto}}1221341\stackrel{{\scriptstyle w_{7}}}{{\mapsto}}1221342$
@@
$1213341\stackrel{{\scriptstyle w_{1}}}{{\mapsto}}1213341\stackrel{{\scriptstyle w_{2}}}{{\mapsto}}1213341\stackrel{{\scriptstyle w_{3}}}{{\mapsto}}1223341\stackrel{{\scriptstyle w_{4}}}{{\mapsto}}1221341\stackrel{{\scriptstyle w_{5}}}{{\mapsto}}1221341\stackrel{{\scriptstyle w_{6}}}{{\mapsto}}1221311\stackrel{{\scriptstyle w_{7}}}{{\mapsto}}1221312$
@@
See Figure 5 for an illustration of Theorem 3.8 for the orbits under the action of $\mathbf{w}$ (i.e., $\pi$ is the identity) over $\mathcal{F}=\operatorname{RG}(5,3)$. This theorem is another instance where there is a homomesy under an action that produces unpredictable orbit sizes, and for which the order of the map is unknown in general.
@@
Figure 5. The two $\mathbf{w}$-orbits for $\mathcal{F}=\operatorname{RG}(5,3)$. The left orbit has length 16 and the right orbit has length 9. Notice that in the left orbit, there are the same numbers (four each) of functions $h$ satisfying any of the conditions $h(2)=1$, $h(3)=1$, $h(4)=1$, or $h(5)=1$. In the right orbit, there are also the same numbers (two each) of functions $h$ satisfying any of the conditions $h(2)=1$, $h(3)=1$, $h(4)=1$, or $h(5)=1$. This illustrates Theorem 3.8 for the case $n=5$, $k=3$, and $\pi$ is the identity.
@@
3.2. Proof of homomesy for whirling restricted growth words
@@
Let $\mathcal{F}$ be either $\operatorname{RG}(n)$ or $\operatorname{RG}(n,k)$ for some $1\leq k\leq n$. Let $\pi,\sigma$ be permutations on $\{2,3,\dots,n\}$ where $\sigma(i)=\pi(i+1)$ for $2\leq i\leq n-1$ and $\sigma(n)=\pi(2)$. Then for any $\mathbf{w}_{\pi}$-orbit $\mathcal{O}=(f_{1},f_{2},\dots,f_{\ell})$, there is a $\mathbf{w}_{\sigma}$-orbit $\mathcal{O}^{\prime}=(f_{1}^{\prime},f_{2}^{\prime},\dots,f_{\ell}^{\prime})$ for which $f_{j}^{\prime}=w_{\pi(2)}(f_{j})$ for all $j\in[\ell]$.
@@
For $h\not=\pi(2)$, $f_{i}(h)=f_{i}^{\prime}(h)$. Also, $f_{i}^{\prime}(\pi(2))=f_{i+1}(\pi(2))$. Thus for any $h\in[n]$, the multiset of values of $f(h)$ is the same for $f$ ranging over $\mathcal{O}$ is the same as that for $\mathcal{O}^{\prime}$. ∎
@@
For example, on $\mathcal{F}=\operatorname{RG}(6)$, $w_{3}\circ w_{5}\circ w_{4}\circ w_{6}\circ w_{2}$ and $w_{4}\circ w_{6}\circ w_{2}\circ w_{3}\circ w_{5}$ satisfy the orbit correspondence above. However, in general two orders for applying the whirling maps do not necessarily yield the same orbit structure if they are not cyclic rotations of each other. For example, $w_{3}\circ w_{5}\circ w_{4}\circ w_{6}\circ w_{2}$ has an orbit structure different from that of $w_{5}\circ w_{6}\circ w_{3}\circ w_{2}\circ w_{4}$. The former has orbit sizes $5,6,6,23,35,62,66$ while the latter has orbit sizes $4,6,20,34,139$.
@@
Let $f\in\mathcal{F}$ where $\mathcal{F}$ is either $\operatorname{RG}(n)$ or $\operatorname{RG}(n,k)$ for some $1\leq k\leq n$. Let $\pi$ be a permutation of $\{2,3,\dots,n\}$. Let $i=\pi(2)$ and suppose $2\leq i\leq n-1$. If $\left(\mathbf{w}_{\pi}^{-1}(f)\right)(i)<\left(\mathbf{w}_{\pi}^{-1}(f)\right)(i+1)$, then $f(i)\not=1$.
@@
[…]
@@
The condition $f(i)=1$ means $1$ and $i$ are in the same block of the partition of $[n]$ corresponding to $f$. Such set partitions can be formed by first choosing a partition of $\{2,3,\dots,n\}$ and placing 1 in the same block as $i$. ∎
@@
Let $\mathcal{O}$ be a $\mathbf{w}_{\pi}$-orbit. The goal is to prove that in $\mathcal{O}$, there are the same number of functions $f\in\mathcal{O}$ satisfying $f(i)=1$ as there are satisfying $f(j)=1$, when $2\leq i,j\leq n$. We will approach this by showing that $\mathcal{O}$ has the same number of functions $f$ satisfying $f(i)=1$ as $f(i+1)=1$, for $2\leq i\leq n-1$.
Let $i\in\{2,3,\dots,n-1\}$. By Remark 3.10, we may assume without loss of generality that $i=\pi(2)$.
@@
By Propositions 3.4 and 3.5, either $(\mathbf{w}_{\pi}(g))(i)=g(i)+1$ or $(\mathbf{w}_{\pi}(g))(i+1)=1$.
@@
By Lemma 3.11, $\left(\mathbf{w}_{\pi}^{m}(f)\right)(i)\not=1$ for all $1\leq m\leq r$. Therefore, given any function $f\in\mathcal{O}$ satisfying $f(i)=1$ and $f(i+1)\not=1$, as we repeatedly apply $\mathbf{w}_{\pi}$ to $f$, we obtain a function that sends $i+1$ to 1 before we get another function that sends $i$ to 1. This implies $\{f\in\mathcal{O}:f(i)=1\}\leq\#\{f\in\mathcal{O}:f(i+1)=1\}$.
By Lemma 3.12, $\{f\in\mathcal{F}:f(i)=1\}=\#\{f\in\mathcal{F}:f(i+1)=1\}$. Therefore, since $\{f\in\mathcal{O}:f(i)=1\}\leq\#\{f\in\mathcal{O}:f(i+1)=1\}$ for every $\mathbf{w}_{\pi}$-orbit $\mathcal{O}$ over $\mathcal{F}$, we must have $\{f\in\mathcal{O}:f(i)=1\}=\#\{f\in\mathcal{O}:f(i+1)=1\}$ for every orbit $\mathcal{O}$. ∎
Theorem 3.8 implies another homomesic statistic for the action of $\mathbf{w}_{\pi}$ on $\operatorname{RG}(n,k)$ only.
@@
Let $\pi$ a permutation of $\{2,3,\dots,n\}$. Under the action of $\mathbf{w}_{\pi}$ on $\operatorname{RG}(n,k)$, the statistic
@@
is homomesic with average k(k2)k(k-2).
@@
Suppose instead that there are $c$ functions $f\in\mathcal{O}$ that satisfy $f(2)=1$. Then this decreases the total value of ${k\choose 2}f(2)$ across $\mathcal{O}$ by ${k\choose 2}c$.
@@
Hence this statistic has average k(k2)k(k-2) on $\mathcal{O}$. ∎
3.3. Whirling on noncrossing RG-words
We finish the paper by considering whirling on the set of RG-words corresponding to noncrossing partitions.